# Quantum enforcement of strong cosmic censorship

### v1

@@ Abstract
We show that the fate of the inner horizon follows directly from the universal structure of quantum field theory in curved spacetime. The same trace anomaly that powers Hawking radiation inevitably destabilizes the Cauchy horizon.
@@
Our analytic result elevates the long-standing Cauchy-horizon instability from a phenomenological observation to a state-independent theorem grounded by QFT in curved spacetime. Unlike previous analyses based on perturbations or numerical simulations of classical mass inflation [14, 4, 5, 6, 15, 16, 17], the present derivation isolates the universal trace anomaly as the dominant driver of curvature blow-up. The Polyakov term alone guarantees the exponential growth $T_{kk}\!\propto\!e^{2\kappa_{-}v}$ and the ensuing null Tipler–strong singularity, ensuring that semiclassical consistency itself enforces the end of predictability. In this precise sense, the trace anomaly provides the fundamental quantum mechanism restoring strong cosmic censorship.
@@ Acknowledgements.
This work is supported by the U.S. Department of Energy, Office of High Energy Physics, under Award Number DE-SC0019470.

### v2

@@ Abstract
We show that the fate of the inner horizon follows directly from the universal structure of quantum field theory in curved spacetime. The same trace anomaly that generates Hawking radiation inevitably destabilizes the Cauchy horizon. A rigorous treatment by Hollands, Wald, and Zahn (HWZ) demonstrated that $\langle T_{\mu\nu}\rangle$ diverges at the Cauchy horizon of Reissner–Nordström–de Sitter black holes using microlocal techniques in four dimensions [8]. Although they noted that the two–dimensional trace–anomaly argument alone was insufficient to capture the full four–dimensional behavior due to mode backscattering, we show that this very simplicity is in fact the key: when expressed in the affine frame, the same trace anomaly fully determines the local divergence mechanism, providing a transparent geometric derivation of the HWZ result and extending it to Kerr, $\Lambda>0$, and nonsingular geometries.
@@
Our analytic result elevates the long-standing Cauchy-horizon instability from a phenomenological observation to a state-independent theorem grounded by QFT in curved spacetime. Unlike previous analyses based on perturbations or numerical simulations of classical mass inflation [15, 4, 5, 6, 16, 17, 18], or the microlocal treatment of HWZ [8], the present derivation isolates the universal trace anomaly as the dominant driver of curvature blow-up. Related semiclassical analyses have also explored anomaly effects on cosmic censorship in lower-dimensional settings, including AdS2 Jackiw–Teitelboim gravity [19] and the rotating BTZ spacetime [20, 21].
The Polyakov term alone guarantees the exponential growth $T_{kk}\!\propto\!e^{2\kappa_{-}v}$ and the ensuing null Tipler–strong singularity, ensuring that semiclassical consistency itself enforces the end of predictability. In this precise sense, the trace anomaly provides the fundamental quantum mechanism restoring strong cosmic censorship.
@@ Acknowledgements.
It is a pleasure to thank Paul Davies, Subhodeep Sarkar and Marija Tomasevic for useful correspondence. This work is supported by the U.S. Department of Energy, Office of High Energy Physics, under Award Number DE-SC0019470.